OpenStudy (anonymous):

Move everything away from y by using the additive inverse/opposite 8x - 3y = 9

OpenStudy (tommo_lcfc):

9 - 3y = 8x 3y = 9 - 8x y = (9-8x)/3

OpenStudy (tommo_lcfc):

or y = 1/3 (9-8x)